# RealCustom++: Representing Images as Real-Word for Real-Time Customization

### v1

@@ Abstract
Text-to-image customization, which takes given texts and images depicting given subjects as inputs, aims to synthesize new images that align with both text semantics and subject appearance. This task provides precise control over details that text alone cannot capture and is fundamental for various real-world applications, garnering significant interest from academia and industry. Existing works follow the pseudo-word paradigm, which involves representing given subjects as pseudo-words and combining them with given texts to collectively guide the generation. However, the inherent conflict and entanglement between the pseudo-words and texts result in a dual-optimum paradox, where subject similarity and text controllability cannot be optimal simultaneously. We propose a novel real-words paradigm termed RealCustom++ that instead represents subjects as non-conflict real words, thereby disentangling subject similarity from text controllability and allowing both to be optimized simultaneously. Specifically, RealCustom++ introduces a novel “train-inference” decoupled framework: (1) During training, RealCustom++ learns the alignment between vision conditions and all real words in the text, ensuring high subject-similarity generation in open domains. This is achieved by the cross-layer cross-scale projector to robustly and finely extract subject features, and a curriculum training recipe that adapts the generated subject to diverse poses and sizes. (2) During inference, leveraging the learned general alignment, an adaptive mask guidance is proposed to only customize the generation of the specific target real word, keeping other subject-irrelevant regions uncontaminated to ensure high text-controllability in real-time. Comprehensive experiments demonstrate the superior real-time customization ability of RealCustom++ in the open domain, outperforming existing state-of-the-art methods in both text controllability and subject similarity. Furthermore, we show that RealCustom++ can be effectively extended to various multiple-subject customization tasks.
@@ 1 Introduction
Figure 1: Illustration of our motivation. (a) Existing paradigm represents the given subject as a pseudo-word ($S^{*}$) and combines it with the given text to collectively guide the generation. The pseudo-word inherently conflicts and entangles with the given text, resulting in the dual-optimum paradox that involves a trade-off between subject-similarity and text-controllability. (b) Our new paradigm, RealCustom++, represents the given subject as a real-word (”toy”) that integrates seamlessly with the given text. The real-word provides accurate influence masks to guide the given subject at each step, ensuring that the subject only influences within the mask while other regions are controlled purely by the given text, thereby achieving disentanglement. (c) The quantitative comparison shows that RealCustom++ achieves superior similarity and controllability to the existing paradigm.
Figure 2: Our RealCustom++ is naturally capable of various customization tasks. (a) One2One: Given a single image depicting the given subject (in open domain, e.g., humans, cartoons, clothes, buildings), RealCustom++ can synthesize images that are consistent with both the semantics of the texts and the appearance of the subjects. (in real-time without any finetuning steps); (b) One2Many: RealCustom++ can decouple and customize each subject in a single reference image. (c) Many2Many: RealCustom++ can customize multiple subjects from multiple reference images. The customized words are highlighted in color.
@@
Text-to-image customization[1, 2, 3], which simultaneously takes given texts and images depicting given subjects as inputs, aims to synthesize new images that are consistent with both the semantics of the texts and the appearance of the subjects. This task allows users to modify specific subjects in provided images using text, such as making your favorite pet wear Iron Man’s suit. Compared to text-to-image generation[4, 5, 6, 7], this task further offers precise control over details that text alone cannot capture (e.g., the unique visual characteristics of your favorite pet), which is crucial for various real-world applications such as film production and personalized recommendations, thereby garnering significant interest from both academia and industry recently. Moreover, this task presents more challenges as its primary goals are dual-faceted: (1) high subject-similarity, i.e., the generated subjects should closely mirror given subjects; (2) high text-controllability, i.e., the remaining subject-irrelevant generated parts should consistently adhere to given texts.
@@
Existing text-to-image customization methods follow a two-step pseudo-word paradigm: (1) representing the given subjects as pseudo-words[2, 3] that share the same dimensionality as real words but are non-existent or meaningless in the vocabulary; (2) combining these pseudo-words with other given texts to guide the generation collectively. According to the different pseudo-words learning ways for subject representation, existing literature can be categorized into finetune-based and encoder-based streams. The former stream[2, 3] adopts the per-subject optimization formulation that finetunes pre-trained text-to-image models[8, 9] to bind the given subject to the pseudo-word. This kind of work either focuses on more efficient finetuning strategies such as finding compact parameter space[10, 11, 12] and extending pseudo-words space[13, 14, 15, 16], or introducing more sophisticated regularization[3, 17, 18] to avoid overfitting. However, this per-subject finetuning requires minutes to hours to learn each subject during inference and over ten reference images containing the same subject to avoid overfitting, which is usually unaffordable for real-world applications. On the contrary, the latter stream[19, 20, 21, 22, 23, 24, 25, 26], which focuses on learning an image encoder to encode the given subject into the pseudo-words, has received more interests since it only requires single forward step and single reference image during inference. These methods focus on developing various adapters[27, 19, 20, 26] to increase the pseudo-word influence for subject-similarity, as well as different regularized losses (e.g., $l_{1}$ penalty [19, 20, 23], alignment loss[25]) to avoid overfitting caused by excessive pseudo-word influence, aiming to maintain a delicate balance. Essentially, existing methods, either the finetune-based or encoder-based, are ensnared in the well-known dual-optimum paradox, where exists a trade-off between subject-similarity and text-controllability and can not achieve the optimum of both simultaneously.
@@
In this study, we argue that the root cause of the dual-optimum paradox lies in the pseudo-word paradigm that integrates pseudo-words representing the given subject with the given text as a unified condition, resulting in inherent conflicts and entanglements between the pseudo-words and the given text. Specifically, (1) conflicts arise from the non-homologous representation between pseudo and real words. Real word representations in pre-trained text-to-image models are learned through large-scale linguistic (e.g., T5[28]) or multi-modal (e.g., CLIP[29]) pre-training, forming semantically consistent contexts. In contrast, pseudo-word representations are learned by reconstructing input subject images (i.e., visual-only), lacking sufficient interaction with real words. Consequently, pseudo-words are typically meaningless and conflict semantically when combined with real words. Increasing the impact of pseudo-words causes real words to deviate from their original semantics, reducing controllability, while decreasing their impact reduces similarity. (2) Entanglements arise from the overlapping influence scope between the given text and subjects. The unified condition guides generation via built-in cross-attention, updating each image region as the weighted sum of all the pseudo and real words, causing given subjects to indiscriminately influence all regions regardless of relevance. Consequently, increasing the impact of pseudo-words in subject-relevant regions to enhance similarity also amplifies their influence in irrelevant regions, making these regions to adhere to the subject rather than the text, thereby reducing controllability, and vice versa. As a typical generation case shown in Figure 1(a), the existing paradigm unifies the pseudo-word (i.e., S representing the given sloth toy) and the given text (i.e., ”in the jungle”) to collectively guide the generation, which results in high subject similarity but low controllability, as the background also closely resembles the input subject images’ background and fails to align with the semantics of ”in the jungle”.
@@
Moreover, the existing pseudo-word paradigm has inherent limitations for open-domain subject customization. It requires explicit correspondence between given subjects and pseudo-words during training, restricting the process to object-level datasets (e.g., OpenImage [30]) with limited categories. Consequently, pseudo-word based methods exhibit poor generalization for real-time, open-domain scenarios.
@@
In this paper, we present RealCustom++, a novel customization paradigm that, for the first time, disentangles subject similarity from text controllability and thereby allows both to be optimized simultaneously without any conflicts. Rather than the existing conflict-entangled paradigm that represents given subjects as pseudo-words, the core idea of RealCustom++ is to represent given subjects as real words (i.e., its super-category word) that can be seamlessly integrated with given texts, and further leveraging the relevance between real words and image regions to disentangle subjects from texts. Specifically, (1) the conflicts can be fundamentally eliminated since these real words are homologous with the given texts, and (2) the superior text-image alignment of pre-trained text-to-image models allows these real words to accurately guide the given subject to influence only subject-relevant regions, thereby eliminating entanglement. As illustrated in Figure 1(b), the specific sloth toy is first represented as the real word ”toy,” which is then integrated with the text ”in the jungle” to form the non-conflict condition ”toy in the jungle.” At each step, this condition generates pure text-guided results with a coherent structure that aim to extract accurate influence mask guidance for the subject. The subject is injected only within the mask, which helps shape a more precise mask for the next step. Through this iterative updating, the generation result of the real word is smoothly and accurately transformed into the subject, while other irrelevant parts are controlled by the text, achieving both high similarity and controllability.
@@
Technically, owing to the absence of annotated paired datasets between ”real-words” and ”subjects”, RealCustom++ introduces an innovative ”train-inference” decoupled framework to realize the novel non-conflict and disentangled real-words paradigm: (1) During training, RealCustom++ learns the alignment between vision conditions and all real words in the text, ensuring high subject-similarity generation in open domains. This is achieved by the Cross-layer Cross-Scale Projector (CCP) to robustly and finely extract subject features, and a Curriculum Training Recipe (CTR) that adapts the generated subject to diverse poses and sizes. (2) During inference, leveraging the learned general alignment, an Adaptive Mask Guidance (AMG) is proposed to only customize the generation of the specific target real word, keeping other subject-irrelevant regions uncontaminated to ensure high text-controllability in real-time.
@@
Specifically, (1) the CCP module functions in three key steps: (i) Initially, the basic features of the subjects are extracted from the last hidden states of pre-trained image encoders, as these states align well with the original text conditions. (ii) These basic features are enhanced for structural robustness by integrating with shallow hidden states via cross-layer attention and enriched with details through integration with higher-resolution hidden states via cross-scale attention. (iii) Finally, the enhanced features are adaptively fused with the original ones without compromising their initial alignment. The final subject features are incorporated into pre-trained diffusion models by extending textual cross-attention with additional visual cross-attention. (2) The Curriculum Training Recipe (CTR) is designed to align the extracted subject features with all real words in the text, facilitating the generation of open-domain subjects in a variety of poses and sizes with harmonious coherence. This is achieved by introducing a novel ”easier-to-harder” subject image data procedure, progressively transitioning the training objective from naive reconstruction to harder re-contextualization. This approach allows the model to converge faster initially and later focus more on aligning with textual semantics, playing a crucial role in bridging the gap between training and inference. (3) During inference, the Adaptive Mask Guidance (AMG) incorporates two distinct branches for each generation step: a Guidance Branch (GuiB) where the visual condition is set to None, and a Generation Branch (GenB) where the visual condition is set to the given subjects. In the GuiB, since cross-attention primarily captures high-level textual semantic correspondences while self-attention captures low-level per-pixel correspondences, we construct the guidance mask by multiplying the low-resolution cross-attention maps of the target native real words with the high-resolution self-attention maps. In the GenB, the guidance mask from GuiB serves as the influence scope for the given subjects, which is multiplied by the outputs of the visual cross-attention to avoid influencing subject-irrelevant parts. Furthermore, considering the guidance mask tends to converge in the middle diffusion steps and become more scattered later, an ”early stop” regularization is proposed for stabilization by reusing previous guidance masks.
@@
We demonstrate that our new paradigm, RealCustom++, is naturally capable of various customization tasks, as illustrated in Figure 2. Given a single image depicting the subject in the open domain, RealCustom++ can generate realistic and harmonious images that consistently adhere to the provided text for the subjects in real-time, without any fine-tuning steps. Moreover, unlike the existing pseudo-word paradigm that learns the correspondence between each pseudo-word and its respective subject, RealCustom++ learns a general alignment between vision conditions and all real words in the text, allowing users to freely select different real words to customize different subjects during inference. Thus, RealCustom++ can naturally extend to multiple subject customizations. For example, RealCustom++ can decouple and customize each subject in a single reference image (i.e., One2Many), as well as customize multiple subjects from multiple reference images (i.e., Many2Many). Both tasks would require additional tedious algorithm design under the existing pseudo-word paradigm.
@@
Conceptual Contributions. For the first time, we (1) identify that the dual-optimum paradox is rooted in the existing pseudo-word paradigm, as it represents subjects as pseudo-words that inherently conflict and entangle with texts; (2) introduce RealCustom++, a novel paradigm that represents subjects as non-conflict real words, thereby disentangling subject similarity from text controllability and allowing both to be optimized simultaneously.
@@
Experimental Contributions. For the first time, we achieve superior subject similarity and text controllability simultaneously for open-domain subjects in real-time scenarios, outperforming existing state-of-the-art customization methods. Moreover, extensive experiments on multiple-subject customization benchmarks demonstrate the exceptional generalizability of our novel paradigm.
@@
Extensions. This manuscript extends RealCustom[31], a conference paper presented at the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) 2024. Building on the same key idea of representing subjects as non-conflict and disentangled real words introduced in our conference version[31], we completely redesign our training framework, enhancing subject similarity by introducing a more structurally robust and detail-rich subject representation and a novel training recipe. In the inference framework, we further redesign the guidance mask construction algorithm, leading to more spatially accurate and temporally stable mask guidance. Moreover, we endow it[31] with the new capability of multiple subject customization. Specifically, the major extensions include:
@@
(1) New Training Framework. (i) New subject representation learning: To enhance subject representation for higher subject similarity, we introduce a novel cross-layer attention mechanism designed to improve structural robustness, along with a cross-scale attention mechanism to enrich detail. These mechanisms are then adaptively integrated, ensuring that the original representation’s alignment with the text is preserved. (ii) New textual alignment learning: A hand-crafted adaptive scoring module was used in our conference version, requiring manually tuned hyper-parameters for optimal generation. We replace it with a novel Curriculum Training Recipe (CTR), which implicitly endows RealCustom++ with the capability to generate given subjects harmoniously in various sizes and poses, thereby eliminating the need for tedious hyper-parameter tuning.
@@
(2) New Guidance Mask Construction Algorithm during Inference. The key to text controllability lies in the quality of the guidance mask during inference. Therefore, we redesign the mask construction: (i) More Spatial Accurate: We introduce a self-attention augmented cross-attention calculation method, which reduces the noise level of naive cross-attention by incorporating per-pixel correlation. (ii) More Temporal Stable: We incorporate an ”early stop” regularization to stabilize the guidance mask in later diffusion steps, which also accelerates the generation speed.
@@
(3) New Capability. (i) Advanced backbones: Previous experiments in our conference version were conducted on SD-v1.5. We further evaluate our paradigm on the more advanced text-to-image backbone model SDXL to thoroughly demonstrate its effectiveness. (ii) Advanced tasks: We introduce a new multi-real-words customization algorithm, thereby extending RealCustom++ to accommodate multiple subject customization across various tasks.

### v2

# RealCustom++: Representing Images as Real Text Word for Real-Time Customization

@@ Abstract
Text-to-image customization, which takes a given text and an image of a given subject as inputs, aims to generate new images by re-contextualizing the subject based on the text. This task allows users to freely modify any specific subject using any textual description, offering more practical applications than text-to-image generation, and has recently garnered increasing interest. Existing works follow the pseudo word paradigm that first represents the subject as a pseudo word (out of the original vocabulary) and then combines it with the given text to generate the image. However, the pseudo word inherently (1) conflicts with the given text because its representation is learned by reconstructing the subject’s visual information, which is completely different from text words, causing the text words to deviate from their semantics when combined, and (2) entangles with the given text due to their overlapping influence scope in each generation step. As a result, the pseudo-word for similarity and the given text for controllability are in opposition, trapping the existing pseudo-word paradigm in a dual-optimum paradox where both cannot be optimal simultaneously. To address this problem, we propose a novel paradigm named RealCustom++ that is completely different from the existing one: at each step, we first combine the subject’s super-category real text word with the text to generate a guidance image with reasonable layout, providing a guidance mask for the subject, and secondly inject the subject only within the mask to help shape a more accurate guidance mask for the next step. Through this iterative updating, regions relevant to the target real text word are accurately shaped into the subject, while other irrelevant ones are guided solely by the text, leading to both high similarity and controllability. Therefore, the conflict will be fundamentally resolved since these real words are homologous with the texts, and the entanglement will be resolved since the guidance mask can accurately disentangle the influence scope for the subject and text. Specifically, RealCustom++ introduces a novel ”train-inference” decoupled framework: (1) During training, we align the visual condition with all words in the text, enabling the model to mirror all visual details of any open-domain subjects. (2) During inference, based on the learned general alignment, we customize the generation of the specific target text word while keeping other subject-irrelevant regions uncontaminated. In contrast to previous methods that excelled in either controllability or similarity, RealCustom++ outperforms them in both, achieving 7.48% and 3.04% improvements in controllability and similarity, respectively. We also achieve a 76.43% improvement in generation quality compared to existing state-of-the-art methods. Furthermore, we propose a novel multiple adaptive mask guidance to extend RealCustom++ to multiple-subject customization, achieving 4.6% and 6.34% improvements in controllability and multiple-subject similarity, respectively. Our work has been commercially applied in JiMeng, and the corresponding link could be found in footnote.
@@ 1 Introduction
Figure 1: Illustration of our motivation. (a) Existing paradigm represents the given subject as a pseudo word ($S^{*}$) and combines it with the given text to generate the image. The pseudo-word inherently conflicts and entangles with the given text, resulting in the dual-optimum paradox that involves a trade-off between subject similarity and text controllability. (b) Our new paradigm, RealCustom++, first represents the given subject as its non-conflicting super-category real word to generate a guidance image with a reasonable layout in the guidance branch, aiming to extract an accurate influence mask for the subject. Then, in the generation branch, the given subject is injected only within the mask while other regions are controlled purely by the given text, thereby achieving disentanglement. See Figure. 3(b) for details. (c) The quantitative comparison shows that RealCustom++ achieves the best similarity and controllability to the existing paradigm simultaneously.
@@
Text-to-image customization[1, 2, 3], which simultaneously takes given texts and images depicting given subjects as inputs, aims to synthesize new images that are consistent with both the semantics of the texts and the appearance of the subjects. This task allows users to modify the specific subject in provided images using any textual description, such as ”making your favorite pet wear Iron Man’s suit”. Compared to text-to-image generation[4, 5, 6, 7], this task further offers precise control over details that text alone cannot capture (e.g., the unique visual characteristics of your favorite pet), which is crucial for various real-world applications such as film production and personalized recommendations, thereby garnering significant interest from both academia and industry recently. Moreover, this task presents more challenges as its primary goals are dual-faceted: (1) high subject-similarity, i.e., the generated subjects should closely mirror given subjects; (2) high text-controllability, i.e., the remaining subject-irrelevant generated parts should consistently adhere to given texts.
@@
Existing text-to-image customization methods follow a two-step pseudo word paradigm: (1) representing the given subject as a pseudo-word[2, 3] that encapsulates the visual subject information, shares the same dimensionality as real words, but is out of the original vocabulary and meaningless; (2) combining the learned pseudo word with other given texts to generate the image. According to the different pseudo words learning ways for subject representation, existing literature can be categorized into finetune-based and encoder-based streams. The former stream[2, 3] learns the pseudo word by finetuning the pre-trained text-to-image models[8, 9] for each subject individually. This kind of work either focuses on more efficient finetuning strategies such as finding compact parameter space[10, 11, 12] and extending pseudo-words space[13, 14, 15, 16], or introducing more sophisticated regularization[3, 17, 18] to avoid overfitting. However, this per-subject finetuning requires minutes to hours to learn each subject during inference and over ten reference images containing the same subject to avoid overfitting, which is usually unaffordable for practical applications. On the contrary, the latter stream[19, 20, 21, 22, 23, 24, 25, 26], which focuses on learning an image encoder to encode the given subject into the pseudo words, has received more interests as it eliminates the cumbersome per-subject finetuning process and requires only a single-step forward inference for all subject once trained. These methods focus on developing various adapters[27, 19, 20, 26] to increase the pseudo word influence for subject similarity, as well as different regularized losses (e.g., $l_{1}$ penalty [19, 20, 23], alignment loss[25]) to avoid overfitting caused by excessive pseudo word influence, aiming to maintain a delicate balance. Essentially, existing methods, whether finetune-based or encoder-based, require a trade-off between subject similarity and text controllability and are therefore ensnared in the well-known dual-optimum paradox that can not achieve the optimum of both simultaneously.
@@
In this study, we argue that the root cause of the dual-optimum paradox lies in the pseudo word paradigm, which integrates pseudo words representing the given subject with the given text as a unified condition to generate the image, resulting in inherent conflicts and entanglements between the pseudo words and the given text. Specifically, (1) conflicts arise from the non-homologous representation between pseudo and real words. Pseudo word representations are learned by reconstructing input subject images in a visual-only manner, lacking context learning with other text words. In contrast, real-word representations are learned through large-scale linguistic context pre-training, such as language models like T5 [28] or multimodal models like CLIP [29], forming semantically consistent contexts. Consequently, these pseudo words typically conflict semantically when combined with real text words because their visual representations are context-independent from other words. Increasing the impact of pseudo-words causes real words to deviate from their original semantics, reducing controllability, while decreasing their impact reduces similarity. (2) Entanglements arise from the overlapping influence scope between the given text and subjects. The combination of the pseudo-word and the given text controls the generation via built-in cross-attention, updating each image region as the weighted sum of all the pseudo and real words. This causes the given subjects to indiscriminately influence all regions, regardless of their relevance. Consequently, increasing the impact of pseudo-words in subject-relevant regions to enhance similarity also amplifies their influence in irrelevant regions, causing these regions to adhere to the subject rather than the text, thereby reducing controllability, and vice versa. As a typical generation case shown in Figure 1(a), the existing paradigm unifies the pseudo-word (i.e., S representing the given sloth toy) and the given text (i.e., ”in the jungle”) to collectively guide the generation. As a result, the background of the generated image closely resembles the input subject image’s background and fails to align with the target background of ”in the jungle” specified in the text, demonstrating high subject similarity but low controllability.
@@
Moreover, the existing pseudo-word paradigm has inherent limitations, as it can only generate well for subjects in seen categories during training, thus restricting its applicability for open-domain subject customization. The primary reason is that the pseudo-word paradigm requires an explicit correspondence between given subjects and pseudo words during training, confining the training process to object-level datasets with limited categories (e.g., OpenImage [30]). Consequently, these pseudo-word-based methods struggle to generate well for subjects in unseen categories, exhibiting poor generalization.
@@
In this paper, we present RealCustom++, a novel customization paradigm that, for the first time, disentangles subject similarity from text controllability and thereby allows both to be optimized simultaneously without any conflicts. The key idea of RealCustom++ is to (1) represent the subject as its non-conflicting super-category real text word to generate a guidance image with a reasonable layout, providing the guidance mask for the subject, and then (2) disentangle the subject from the text by ensuring that the subject only influences the subject-relevant region within the guidance mask while other regions are purely controlled by the given text. Therefore, (1) conflicts can be fundamentally eliminated since these real words are homologous with the given texts, and (2) entanglement can be eliminated since the guidance mask can accurately disentangle the influence scope for the given subject and given text. As illustrated in Figure 1(b), the specific sloth toy is first represented as the real word ”toy”, which is then integrated with the text ”in the jungle” to form the non-conflicting control condition ”toy in the jungle”. At each step, this control condition is used in the guidance branch to generate the pure text-guided result with a reason layout and structure, which aims to extract accurate influence mask guidance for the subject. Then in the generation branch, the given subject is injected only within the mask, which helps the guidance branch to shape a more precise mask in the next step. Through this iterative updating, the generation result of the real-word is smoothly and accurately transformed into the subject, while other irrelevant parts are controlled by the text, achieving both high similarity and controllability.
@@
Technically, owing to the absence of annotated paired datasets between ”real words” and ”subjects”, RealCustom++ introduces an innovative ”train-inference” decoupled framework to realize the novel non-conflicting and disentangled real words paradigm: (1) During training, RealCustom++ learns the alignment between vision conditions and all real words in the text, ensuring high subject similarity generation in open domains. This is achieved by the proposed novel Cross-layer Cross-scale Projector (CCP) to robustly and finely extract subject features, and the proposed novel Curriculum Training Recipe (CTR) that adapts the generated subject to diverse poses and sizes. (2) During inference, leveraging the learned general alignment, we propose the Adaptive Mask Guidance (AMG) to only customize the generation of the specific target real word, keeping other subject-irrelevant regions uncontaminated to ensure high text-controllability in real-time.
@@
Specifically, (1) the CCP module operates in three key steps: (i) Initially, the basic features of the subjects are extracted from the last hidden states of pre-trained image encoders, as these states align well with the original text conditions. (ii) These basic features are enhanced for structural robustness by integrating with shallow hidden states via cross-layer attention and enriched with details through integration with higher-resolution hidden states via cross-scale attention. (iii) Finally, the enhanced features are adaptively fused with the original ones without compromising their initial alignment. The final subject features are incorporated into pre-trained diffusion models by extending textual cross-attention with additional visual cross-attention. (2) The Curriculum Training Recipe (CTR) is designed to align the extracted subject features with all real words in the text, facilitating the generation of open-domain subjects in a variety of poses and sizes with harmonious coherence. This is achieved by introducing a novel ”easier-to-harder” subject image data procedure, progressively transitioning the training objective from naive reconstruction to harder re-contextualization. This approach allows the model to converge faster initially and later focus more on aligning with textual semantics, playing a crucial role in bridging the gap between training and inference. (3) During inference, the Adaptive Mask Guidance (AMG) incorporates two distinct branches for each generation step: a Guidance Branch where the visual condition is set to None, and a Generation Branch where the visual condition is set to the given subjects. In the Guidance Branch, since cross-attention primarily captures high-level textual semantic correspondences while self-attention captures low-level per-pixel correspondences, we construct the guidance mask by multiplying the low-resolution cross-attention maps of the target native real words with the high-resolution self-attention maps. In the Generation Branch, the guidance mask from Guidance Branch serves as the influence scope for the given subjects, which is multiplied by the outputs of the visual cross-attention to avoid influencing subject-irrelevant parts. Furthermore, considering the guidance mask tends to converge in the middle diffusion steps and become more scattered later, an ”early stop” regularization is proposed for stabilization by reusing previous guidance masks.
@@
We demonstrate that our new paradigm, RealCustom++, addresses the inherent defects of the previous pseudo word paradigm in not generalizing to unseen categories and is capable of various open-domain customization tasks, as illustrated in Figure 2. Given a single image depicting the subject in the open domain, RealCustom++ can generate realistic and harmonious images that consistently adhere to the provided text for the subjects in real-time, without any fine-tuning steps. In contrast to the existing pseudo-word paradigm that learns the correspondence between each pseudo-word and its respective subject, RealCustom++ learns the general alignment between vision conditions and all real words in the text during training, allowing users to freely select any real words to customize any subjects during inference. Therefore, RealCustom++ can customize subjects of any category during inference and exhibit strong generalization for real-time, open-domain scenarios. Furthermore, RealCustom++ can naturally extend to multiple subject customizations by selecting multiple real words together. For example, RealCustom++ can decouple and customize each subject in a single reference image (i.e., One2Many), as well as customize multiple subjects from multiple reference images (i.e., Many2Many). Both tasks would require additional tedious algorithm design under the existing pseudo-word paradigm.
@@
Conceptual Contributions. We (1) identify that the dual-optimum paradox is rooted in the existing pseudo-word paradigm, as it represents subjects as pseudo-words that inherently conflict and entangle with texts; (2) introduce RealCustom++, a novel paradigm that represents subjects as non-conflicting real words, thereby disentangling subject similarity from text controllability and allowing both to be optimized simultaneously.
@@
(iv) Moreover, we further propose a novel multiple adaptive mask guidance to effectively extend RealCustom++ to various multiple-subject customization tasks.
@@
Experimental Contributions. To the best of our knowledge, for the first time, we achieve superior controllability and similarity simultaneously for open-domain subjects in real-time scenarios, outperforming the previous best method in controllability by 7.48% on SDXL measured by CLIP ViT-L/14 textual score, the previous best method in similarity by 3.04% on SDXL measured by CLIP ViT-L/14 visual score, and previous best method in image quality by 76.43% measure by Image Reward score, respectively. Moreover, as for the multiple subject customization, we achieve 4.6% and 6.34% improvement in text controllability and multiple-subject similarity.
@@
Extensions. This manuscript extends RealCustom[31], a conference paper presented at the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) 2024. Building on the same motivation, we completely redesign the two key algorithm and hence rebuild our paradigm, i.e., (1) Training Paradigm Redesign: We propose a novel cross-layer cross-scale projector to extract subject features in a robust and rich-detailed manner, and a novel curriculum training recipe to adapt the generate subject to diverse pose and size, significantly enhancing the subject similarity; (2) Guidance Mask Algorithm Redesign: During inference, we redesign the proposed adaptive mask guidance, leading to more spatially accurate and temporally stable mask guidance for better text controllability. Moreover, we endow it[31] with the new capability of multiple subject customization. Finally, we achieve a significant improvement over conference version on both similarity and controllability.
Specifically, the major extensions include:
@@
(1) New Training Algorithm. (i) New subject representation learning: To enhance subject representation for higher subject similarity, we introduce a novel cross-layer attention mechanism designed to improve structural robustness, along with a cross-scale attention mechanism to enrich detail. These mechanisms are then adaptively integrated, ensuring that the original representation’s alignment with the text is preserved. (ii) New textual alignment learning: A hand-crafted adaptive scoring module was used in our conference version, requiring manually tuned hyper-parameters for optimal generation. We replace it with a novel Curriculum Training Recipe (CTR), which implicitly endows RealCustom++ with the capability to generate given subjects harmoniously in various sizes and poses, thereby eliminating the need for tedious hyper-parameter tuning.
@@
(2) New Guidance Mask Algorithm. The key to text controllability lies in the quality of the guidance mask during inference. Therefore, we redesign the mask construction: (i) More Spatial Accurate: We introduce a self-attention augmented cross-attention calculation method, which reduces the noise of naive cross-attention by incorporating per-pixel correlation. (ii) More Temporal Stable: We propose an ”early stop” regularization to stabilize the guidance mask in later diffusion steps, which also accelerates the generation.
@@
(3) New Capability. (i) Advanced backbones: Previous experiments in our conference version were conducted on SD-v1.5. We further evaluate our paradigm on the more advanced text-to-image backbone model SDXL to thoroughly demonstrate its effectiveness. (ii) Advanced tasks: We introduce a novel multiple adaptive mask guidance algorithm, thereby extending RealCustom++ to accommodate multiple subject customization across various tasks.
(4) Higher Performance. Compared to our conference version, we achieve improvements of 6.06% in text controllability and 2.32% in subject similarity simultaneously, and a significant 56.6% improvement on generation quality.
Figure 2: Our RealCustom++ is capable of various customization tasks. (a) One2One: Given a single image depicting the given subject (in open domain, e.g., humans, cartoons, clothes, buildings), RealCustom++ can synthesize images that are consistent with both the semantics of the texts and the appearance of the subjects. (in real-time without any finetuning steps); (b) One2Many: RealCustom++ can decouple and customize each subject in a single reference image. (c) Many2Many: RealCustom++ can customize multiple subjects from multiple reference images. The customized words are highlighted in color.